# SMILES & SELFIES have to go: Representation of Molecules via Algebraic Data Types

### v2

@@ 3.13 Fingerprints
Fingerprints involve encoding key features of a molecule, typically into a vector representation and can offer convenience for machine learning applications. However, a significant drawback of fingerprints is their inability to consistently represent tautomers, which are molecules that have the same atomic composition but different structural arrangements. Tautomers often have distinct molecular structures, leading to different encodings in the bitmaps or fingerprints used for molecular database searching [34]. This inconsistency can result in the failure to identify the presence of a particular molecule in a database accurately.
@@
A fundamental issue with numeric vector representations, such as fingerprints, is the lack of inherent semantics associated with the individual numbers or floating-point values. Without additional contextual information or a predefined mapping, the numeric values in the vector do not carry intrinsic meaning. This lack of semantics may hinder the interpretability and reasoning capabilities of machine learning models trained on such representations.
In contrast, algebraic data types (ADTs) provide a structured and type-safe representation that inherently encodes the semantics of the molecular data. By defining the data structure and its associated types, ADTs constrain the space of possible values and operations, ensuring that only meaningful and valid transformations can be applied. This type-driven approach helps prevent errors and inconsistencies that may arise from arbitrary manipulation of numeric vectors.
Furthermore, the type information associated with ADTs can guide and constrain the machine learning and software development process [^3] See Conor Mcbride talk about this issue: https://www.youtube.com/watch?v=v2yXrOkzt5w. The type constraints restrict the search space and focus learning algorithms on meaningful and plausible solutions. This may well lead to more efficient and effective learning, as the model can leverage the structure and semantics embedded in the ADT representation.
@@
Fingerprints are of course useful, but can be defined on the more general type data type by the programmer. For instance, given MAP4 touted as “One Molecular Fingerprint to rule them all” [35], can be constructed from an ADT, but fingerprints can lose information, cause bit collisions and lose algebraic properties of the elements of the fingerprint. These algebraic properties such as group theoretic properties can be used to predict molecular properties. Variational Auto Encoders (VAEs) also have problems, [6], whom claim that obtaining structured output to be a hard and currently unsolved problem, even for a simple graph data structure.
3.14 RDKit as representation
One might also ask whether existing and widely used Cheminformatics libraries support representation. Behemoth Cheminformatics programming libraries, such as RDKit, [10] and lightweight molecular manipulation libraries that exist on top of these e.g. datamol [^4] https://datamol.io/ and https://www.rdkit.org/ (which builds on NumPY [36], Pandas, RDKit and Python 3+), Datamol explicitly [^5] https://doc.datamol.io/stable/tutorials/Preprocessing.html don’t advocate themselves as providing a prototypical representation. These libraries parse representations to computational representations and then apply functions to them, before outputting them to strings like SMILES.
In the RDKit’s Mol classes, molecules are stored as a graph in the form of an adjacency list, with optional extra information in the form of properties and ring information. The internal representation therefore inherits the limitations of 2D graph based representations already discussed for computational purposes. Bonds in the graph can be labeled with 22 pre-defined bond types, and limited stereochemical features can be described, again with labeling.
The computational representation of RDKit in C++ is complex, taking up over 400 lines of highly complex code and the molecule class taking up 873 lines of code as of 2022. There are no clear invariants, or guarantees about the management of side effects in snippets shown below. The ubiquitous presence of null pointers and highly abstruse C++ keywords, scattered throughout the code, insinuates points of failure of the program. It is unclear, whether and when the program may crash and what the consequences might be. The following snapshot of lines $792$ to $821$ in ROMol.h [^6] https://www.rdkit.org/docs/cppapi/ROMol_8h_source.html which represents the molecule class, provides a brief glimpse into the extreme complexity of code:
1 #ifdef RDK_USE_BOOST_SERIALIZATION
[…]
3 //! @{
4 template <class Archive>
5 void save(Archive &ar, const unsigned int version) const;
6 template <class Archive>
7 void load(Archive &ar, const unsigned int version);
8 BOOST_SERIALIZATION_SPLIT_MEMBER()
9 //! @}
10 #endif
@@ 8 BOOST_SERIALIZATION_SPLIT_MEMBER()
12 private:
13 MolGraph d_graph;
14 ATOM_BOOKMARK_MAP d_atomBookmarks;
15 BOND_BOOKMARK_MAP d_bondBookmarks;
16 RingInfo *dp_ringInfo = nullptr;
17 CONF_SPTR_LIST d_confs;
18 std::vector<SubstanceGroup> d_sgroups;
19 std::vector<StereoGroup> d_stereo_groups;
20 std::unique_ptr<boost::dynamic_bitset<>> dp_delAtoms = nullptr;
21 std::unique_ptr<boost::dynamic_bitset<>> dp_delBonds = nullptr;
22
23 friend RDKIT_GRAPHMOL_EXPORT std::vector<SubstanceGroup> &getSubstanceGroups(
24 ROMol &);
25 friend RDKIT_GRAPHMOL_EXPORT const std::vector<SubstanceGroup>
26 &getSubstanceGroups(const ROMol &);
27 void clearSubstanceGroups() { d_sgroups.clear(); }
28
29 protected:
30 unsigned int numBonds{0};
Listing 1: This is just a small fragment of the molecule class in RDKit. It demonstrates the sheer complexity and obfuscation of a concept that can be represented more simply in the ADT proposed in this article.
We aim to demonstrate that our ADT-based representation of molecules in Haskell offers a more concise, expressive, and type-safe approach. By leveraging the power of algebraic data types and the strong static typing of Haskell, we can define molecular structures in a declarative and compositional manner. The resulting code is more readable, maintainable, and less prone to runtime errors. The use of pure functions and the explicit management of side effects in a pure functional language such as Haskell, will enable easier reasoning about the behavior of the program and facilitate formal verification techniques.
Moreover, we claim that the ADT-based representation provides a clear separation of concerns between the data structure and the operations performed on it, interface versus implementation. This allows for a separation of concerns between those who are designing better algorithms for the data type and those who only need to use the data type. The use of higher-order functions and type classes in Haskell further enhances the reusability and composability of the code.
While cheminformatics libraries like RDKit and datamol have been widely adopted and have demonstrated their utility in various applications, their complex often OOP-based representations can hinder the development and maintenance of robust and reliable software. The proposed ADT-based approach offers a compelling alternative that prioritizes simplicity, type safety, and functional mathematical rigor.
@@ 3.15 Molecules and programming languages
This makes the proposal here the first to suggest to cheminformaticians to use constructs rooted in functional programming languages to be directly involved as representation. The notion that “molecular programming languages” could be possible, occurred in an article about the future of SELFIES [1]. However, the sense attached to that project has been questioned previous to this.
Investigations of type safe functional programming for cheminformatics has occurred before. Recent work has suggested the use of OCaml in writing software for cheminformatics [39, 40]. However, the emphasis in both of these articles as well as the Haskell library Ouch [41], was not on molecular representations but more a tutorial on functional programming. This is highlighted by the papers introduction of lengthy tutorials on functional programming. Furthermore they do not show how Bayesian Inference can be performed on drug structures over algebraic data types which is essential in motivating efficiency improvements in not requiring exploration of syntactically nonsensical regions of molecular space.
4 Probabilistic programming
As discussed in a recent thesis on modular probabilistic programming, and I quote from it through this and the next paragraph [42], the Bayesian approach to probabilistic reasoning is known as Bayesian Inference, and is an increasingly popular method to perform machine learning that enjoys strong mathematical foundations. Bayesian Inference has useful properties, compared to deep learning based machine learning, such as interpretability, explainability and expressiveness with respect to various forms of uncertainty [43, 44]. At the heart of the Bayesian paradigm is the notion of a generative model. Generative models can account for the expressiveness of uncertainty and interpretability. In order to perform probabilistic modelling, knowledge about a domain and the uncertainty associated with that knowledge is encoded into a generative model, using expert knowledge. Inference algorithms estimate parameters of the model, that generate the observed data, by conditioning models on associated observed data. For instance, a climate researcher may create a basic model of a city when studying urban heat islands. The climate researcher would condition the model of the city with empirical data, such as the observed temperatures of the city, in order to infer values of interest such as latent city temperatures or factors that influence the temperature of a city. Probabilistic programming languages provide a generic framework to perform inference.
To highlight the importance of probabilistic programming languages, large multinational technology companies such as Microsoft, Google and Uber have recently started to invest in their own probabilistic programming languages (Infer.NET [45], Edward [46] and Pyro [47] respectively). Increasingly, machine learning use cases find their way into areas as diverse as critical defense national infrastructure [48][49] or medical diagnostics [50]. Society is therefore increasingly in a position that relies on programmers to produce correct machine learning related code. As probabilistic programming becomes more important, the code that underlies probabilistic programming languages will face increasing scrutiny. Fortunately, there are two important programming language research goals which aim to ensure the correctness of code; modularity and type-safety. In particular, they ensure that programmers can be certain early on – at compile time – that composing code will not cause programs to crash or introduce unintended (inference) algorithmic bias. These research goals have two effects; as programming libraries become safer and more modular, machine learning models will become cleaner, simpler and easier to maintain, which will additionally make machine learning more trusted, and perhaps more widely adopted.
@@ 4 Probabilistic programming
In the context of molecular representation, probabilistic programming offers a natural way to capture the inherent uncertainty and variability associated with molecular structures and properties. By combining the expressive power of functional programming with the ability to represent and reason about probability distributions, PPLs enable the development of rich, probabilistic models of molecules.
5 Methodology: Laziness and The Algebraic Data Type.
Our molecular Algebraic Data Type is written in Haskell[53], a lazy, purely functional, type-safe language, with type classes and algebraic data types.
@@
Thus, if you only access the first 10 molecules in the dataset, only 10 calculations take place; the remaining 999,990 molecular properties remain as unevaluated thunks.

### v3

@@ 1.15 Molecular Fingerprints
There are also drawbacks to fingerprints such as their inability to consistently represent tautomers. Tautomers often have distinct molecular structures, leading to different encodings in the bitmaps or fingerprints used for molecular database searching [38]. This inconsistency can result in the failure to identify the presence of a particular molecule in a database accurately.
@@
A potential issue with purely floating point or integer vectors, such as fingerprints, is the lack of corresponding semantics in the molecular world. Without additional contextual information or a predefined mapping, the numeric values may represent invalid or physically impossible molecules. This may well hinder the interpretability and reasoning capabilities of machine learning models trained on such representations.
@@
Fingerprints are of course useful, but can be defined on the more general type data type by the programmer. For instance, given MAP4 touted as “One Molecular Fingerprint to rule them all” [35], can be constructed from an ADT, but fingerprints can lose information, cause bit collisions and lose algebraic properties of the elements of the fingerprint.
1.16 RDKit as a Representation
RDKit is widely used in cheminformatics for molecular manipulation, property prediction, and conversion between molecular representations. However, its primary function is not to serve as a standalone molecular representation but rather to process, transform, and analyze existing representations such as SMILES strings and MOL/SDF files.
Internally, RDKit uses the Adjacency List from the Boost Graph Library to represent molecules as 2D graphs with additional metadata. Bonds are extensively labeled with predefined bond types, while stereochemical features are captured through limited labeling schemes. This approach inherits the fundamental weaknesses of graph-based representations previously discussed, including ambiguity in encoding stereochemistry and the necessity of extensive labeling to capture molecular nuances, although 3D information may be stored in the “Conformer” object.
While RDKit’s internal representation is based on adjacency lists, we refrain from formally describing its data type as such, as we lack proof of its correctness. The complexity of RDKit’s implementation makes it difficult to verify whether the adjacency list truly functions as intended. Moreover, as demonstrated in Listing 1 and 2, mutable state introduces the potential for undefined behavior, reinforcing concerns about the safety and correctness of the representation.
RDKit’s reliance on a non-functional paradigm means that operations on molecular graphs can lead to unintended side effects, making the system more prone to errors, although we acknowledge this criticism is not unique to RDKit, and is part of a broader criticism of the non-functional programming language paradigm.
The reliance on labeling in RDKit exemplifies the inherent issues of such an approach. For example, a bond order of 2.5 is labeled as TWOANDAHALF, an inelegant solution reminiscent of criticisms raised by Dietz on the limitations of labeled representations. Instead of a structural or algebraic description of bond order, RDKit resorts to a categorical label, which may introduce inefficiencies and inconsistencies when interfacing with other systems.
Datamol, a higher-level library built atop RDKit, further highlights the limitations of RDKit as a representational system. Primarily designed for molecular conversion and computation, Datamol features 45 “featurizers”, all of which take SMILES strings as input. The framework is not optimized for storage or transmission but rather for extracting graph-based molecular features for machine learning applications. This dependence on SMILES as an intermediary highlights a fundamental issue: Datamol and similar tools do not provide an independent molecular representation but instead rely on conversion mechanisms.
1Some care must be taken when changing the structure of a graph (via adding or removing edges). Depending on the type of adjacency_list and on the operation, some of the iterator or descriptor objects that point into the graph may become invalid. For example, the following code will result in undefined (bad) behavior:
2
3typedef adjacency_list<listS, vecS> Graph; // VertexList=vecS
4Graph G(N);
5// Fill in the graph...
6
7// Attempt to remove all the vertices. Wrong!
8graph_traits<Graph>::vertex_iterator vi, vi_end;
9for (tie(vi, vi_end) = vertices(G); vi != vi_end; ++vi)
10 remove_vertex(*vi, G);
@@ 1.16 RDKit as a Representation
12// Remove all the vertices. This is still wrong!
13graph_traits<Graph>::vertex_iterator vi, vi_end, next;
14tie(vi, vi_end) = vertices(G);
15for (next = vi; vi != vi_end; vi = next) {
16 ++next;
17 remove_vertex(*vi, G);
18}
Listing 1: Example description of undefined behaviour which can cause errors, taken from the Boost GraphLibrary documentation, illustraing nature of data type and operations which can be performed as of Wed 5 Feb 2025.
1// ROFIX: these will need to be readonly somehow?
2typedef std::map<int, ATOM_PTR_LIST> ATOM_BOOKMARK_MAP;
3typedef std::map<int, BOND_PTR_LIST> BOND_BOOKMARK_MAP;
Listing 2: Example listing as of Wed 5 Feb 2025.
Moreover, RDKit’s use of null pointers introduces potential points of failure, reflecting broader concerns about type safety in non-functional programming paradigms. The lack of strong guarantees about memory safety and the absence of robust type constraints can lead to errors that are difficult to detect and debug. While this criticism extends beyond RDKit to imperative and object-oriented programming more generally, it underscores the representational fragility of systems that rely on mutable state and loosely enforced data structures.
@@ 1.17 Molecules and programming languages
1.18 Functional languages and cheminformatics
Work using functional programming in cheminformatics has been done before, however the focus of such efforts were usually related to ease of writing or safety of code, or efficiency of functional programming languages, rather than attempting to introduce a new representation by explicitly considering data types.
Ouch [41] is a chemical informatics toolkit written entirely in Haskell, released in 2010. It was released “in the hope that it will be useful”, but has seen little development or since 2013. Internally, molecules are represented with a molecular graph, implemented as an algebraic data type, where bond types are categorically labeled (e.g. “sigma”, “pi”, or “aromatic”), and without any support for stereochemical features or 3D information.
Chemf[10] is a toolkit written in Scala. Its purpose was not to create a molecular representation, but to permit existing cheminformatics workflows to utilise the functional programming paradigm, and to avoid the problems which can occur with mutable state and null pointers. The representation used is a 2D molecular graph written as a connectivity list.
“Radium” is partially complete Haskell library. It has support for atomic orbital information, but molecules may only be encoded as a linear notation based upon SMILES strings. Interestingly, the data types used for molecular representation are algebraic, and while being based upon SMILES syntax, are not in fact “strings”.
Other work [9] has suggested using OCaml for cheminformatics, however this work was primarily written as an introduction to chemists for functional programming, with a granular explanation of specific functions, and a focus on ease of writing code, type safety, and efficiency, such as the ease of parallelization.
This makes the proposal here the first to suggest to cheminformaticians use constructs rooted in functional programming languages directly in the representation of molecules themselves. Additionally, unlike previous work in functional programming, we attempt to expand the scope of molecular representation, and explicitly consider the role of the data type and how this affects machine learning, including with concepts such as Bayesian machine learning and Geometric deep learning.
1.19 Geometric Deep Learning on Molecular Representations
Geometric deep learning (GDL) extends deep learning techniques to non-Euclidean domains such as graphs, manifolds, and meshes[2]. In the context of molecular modeling, molecules can be represented as graphs, where atoms serve as nodes and chemical bonds as edges. This formulation allows GDL methods, particularly graph neural networks (GNNs), to learn representations that capture topological and geometric invariances or equivariances of molecules [12, 13].
A key advantage of GDL in cheminformatics is its ability to incorporate geometric priors, such as rotational and permutation invariance, ensuring that appropriate properties are invariant to the molecular representation at hand, in particular the ID of an Atom (by leveraging permutation invariance). These methods have already been applied successfully in deciphering interaction fingerprints from protein molecular surfaces [42].
The application of GDL is particularly relevant in addressing the limitations of traditional string-based molecular representations, such as SMILES, which encode chemical connectivity but neglect spatial conformation and stereochemistry [1]. By directly operating on molecular graphs and 3D structures, GDL-based approaches provide a more expressive and chemically meaningful representation, enhancing predictive performance in downstream tasks [13].
Our representation of the bond matrix, is indeed a graph structure, computationally implemented with a Haskell Map for the sake of speed. Thus, graph algorithms from the field of geometric deep learning can be quickly and simply applied to our representation. Given these advantages, GDL presents a compelling framework for molecular representation, motivating its adoption as an alternative to conventional symbolic encodings in cheminformatics.
1.20 Probabilistic programming
Typically, in conventional programming, one provides inputs to a function to compute outputs. However, in Bayesian inference, the process is reversed: given observed outputs, we infer the likely inputs and function parameters that could have produced them. For instance, given a model that outputs the solubility, for which we have the values, the inference process determines the parameter values that generated those data given in the training set, with the aim of finding parameter values that generalise the underlying data distribution. In the context of computational chemistry, the input parameters could represent any variable (parameter) that determines the property of a molecule, in a particular representation. This means the feasibility of sampling is fundamentally constrained by the operations permitted on the type of data at hand. If this was used with a String datatype, the notion of sampling and scoring becomes.
This approach, is known as Bayesian inference and is automatically performed by probabilistic programming languages. It is an increasingly popular method to perform machine learning that enjoys strong mathematical foundations [43]. Large technology companies such as Uber, Google and Microsoft, have invested in their own probabilistic programming languages (Pyro [44] Infer.NET [45], Edward [46],respectively). Academia has also focussed on building modular probabilistic programming languages which allow composing inference algorithms to build other inference algorithms [47, 48, 49].
Bayesian Inference has useful properties, compared to deep learning based machine learning, such as interpretability, explainability and expressiveness with respect to various forms of uncertainty [50, 51]. At the heart of the Bayesian paradigm is the notion of a generative model. Generative models can account for the expressiveness of uncertainty and interpretability. Probabilistic programming languages provide a generic framework to perform inference. Bayesian inference has already been used in drug design e.g. [52], where a Bayesian neural network is used to predict drug toxicity.
A Bayesian generative model is defined as a program that at least contains the following two operations:
  • The sample effect (or operation), equips models with an operation that returns a uniformly distributed probability floating point value from the unit interval, expressing probabilistic non-determinism. This is a natural operation in many areas of science which use probability density functions such as sampling from an Orbital or the radioactive decay of a particle (subject to building an appropriate distribution).
  • The score effect (or operation), provides a mechanism to calculate the likelihood of model samples, given observed data.
There are three well-established methods of Bayesian inference:
  • Monte-Carlo Inference – Uses sampling-based methods like MCMC to approximate posterior distributions in probabilistic programs.
  • Variational Inference – Converts inference into an optimization problem by approximating the posterior with a simpler distribution.
  • Exact Inference – Computes the posterior distribution analytically when closed-form solutions exist, often feasible only for simple models.
In this article, for the sake of simplicity, we will use Trace-Metropolis Hastings, although the example below uses rejection sampling which is another Monte-Carlo method. Rejection sampling becomes inefficient in high-dimensional spaces, where too many samples get rejected. In Trace-Metropolis Hastings, a sequence of random samples is generated by iteratively proposing new parameter values and accepting or rejecting them based on a probability ratio. This process explores the posterior distribution by permuting the sampled values to refine the parameter estimates that best generate the observed data.
@@ 1.20 Probabilistic programming
In the context of molecular representation, probabilistic programming offers a natural way to capture the inherent uncertainty and variability associated with molecular structures and properties. By combining the expressive power of functional programming with the ability to represent and reason about probability distributions, PPLs enable the development of rich, probabilistic models of molecules. Consider a simple example involving inferring the bias of a coin in webPPL [53].
Example of inferring the bias of a coin with probabilistic programming:
1 // Inferring the bias of a coin using rejection sampling.
2 var heads = true, tails = false;
3 var observed_data = [heads, tails, heads, heads]
4
5 var model = function() {
6 var prior = sample(Uniform({a: 0, b: 1}))
7 var likelihood = function(p) { return bernoulli(p); };
8 var samples = mapData({data: [prior, prior, prior, prior]}, likelihood);
9 condition(samples.toString() $=$ observed_data.toString())
10 return prior
11 }
12
13 // Plot the result of inference.
14 viz(Infer({method: rejection’, samples:1000}, model))
Listing 3: Inferring the bias of a coin using rejection sampling in WebPPL
Figure 2: WebPPL inference graph of biased coin inference, based on the code. The most likely bias parameter is 0.75, which makes sense as 3/4 of the coins are heads (true).
The prior is sampled from the uniform interval $[0,1]$ and the likelihood function is Bernoulli, assuming values in the set $\{0,1\}$. The conditioning process is explicit with condition. Conditioning is the fundamental method in which statistical inference over parameters is done [54]. Only program traces that fulfill the condition are considered by the infer method, where a trace, defined by [54], is simply the tree data structure that captures the random choices encountered and the path taken by the interpreter while evaluating the program. The condition operation is a specialization of score to be either exactly the observed data or not. Some scoring functions use distance measures.
The random primitives are in this case real-valued random variables and Boolean random variables. The method of inference used here is a non-Markov Monte Carlo approach, termed rejection sampling. Intuitively, this can be understood by considering the traces of the program that progress past the condition statement, which will generally be dominated by prior values that tend to yield the observed data. Based on observed data, one could also include knowledge about scientific laws restricting the world, resulting in $\theta$ values that are most “realistic”.
In [55, 56], there are discussions of the importance of classifiers providing not just a single output, but two key pieces of information: a class label and an associated probability. This dual output allows for a more nuanced understanding of the classifier’s confidence in its predictions. For instance, in binary classification tasks, a classifier might assign a label (e.g., ‘spam’ or ‘not spam’) and also provide the probability of the instance belonging to that class. This probabilistic information is crucial for tasks such as ranking, where items are ordered based on the likelihood of belonging to a particular class, and for making informed decisions in applications where the costs of false positives and false negatives differ. [56] emphasizes that effective classification involves not only assigning the correct label but also accurately estimating the probability associated with that label to reflect the classifier’s certainty. In the case of SMILES, an estimate of 0.5 that it the molecule is soluble may mean exactly that or it may be saying that it is an example that is has never seen before. Without a grammar that tells one what is valid and what is invalid, it relegates the probability estimates to be ambiguous and potentially meaningless.
1.21 Laziness
@@
Thus, if you only access the first 10 molecules in the dataset, only 10 calculations take place; the remaining 999,990… molecular properties remain as unevaluated thunks.
@@
2 Methodology
2.1 Our Algebraic Data Type
Definition 9 (Algebraic Data Type).
An algebraic data type (ADT) over a set of base types is any type built up from:
  1. 1.
    Sums (coproducts): If $A$ and $B$ are types, then $A+B$ is the type of “either an $A$ or a $B$”;
  2. 2.
    Products (Cartesian products): If $A$ and $B$ are types, then $A\times B$ is the type of pairs $(a,b)$ with $a\in A$ and $b\in B$;
  3. 3.
    Recursion via a least fixpoint: Suppose $F$ is an operation that takes a type and returns a new type. The type $\mu X.\,F(X)$ is then the smallest (or simplest) solution to the equation $T\cong F(T)$. Intuitively, $T$ is built exactly in the shape that $F$ prescribes, by referencing itself over and over as needed.
    This construction gives a natural way to represent (iso)recursive data structures. Because $T$ can contain smaller copies of itself, you can define objects like lists, trees, and other self-referential shapes. Calling $\mu X.\,F(X)$ the “least fixpoint” emphasizes that it includes no extra elements beyond what arise from repeatedly applying $F$. In practice, this ensures that the type exactly captures the recursive shape you intend, without unnecessary complexity.
Concretely, any recursive ADT can be expressed as
$T\;=\;\mu X.\,F(X),$
where $F$ uses only sums, products, and previously defined base types.
For those, familiar with category theory, $\mu X.\,F(X)$ is the initial $F$-algebra, ensuring that it is the “smallest” type satisfying $T\cong F(T)$. Equirecursive type theories treat $\mu X.\,F(X)$ and $F(\mu X.\,F(X))$ as definitionally equal, while isorecursive theories regard them only as isomorphic via explicit fold/unfold operations.
Our molecular Algebraic Data Type is written in Haskell [59], a lazy, purely functional, type-safe language, with type classes and algebraic data types. In languages like Haskell, the recursion is effectively isorecursive: data definitions (e.g. data List a = Nil | Cons a (List a)) introduce an implicit fold/unfold mechanism (via constructors and pattern matching) even though it appears syntactically direct.